whaat do you know about the the value of x or the value of y if xy=0
Answer by haileytucki(390)

xy=0
Divide each term in the equation by y.
(xy)/(y)=(0)/(y)
Cancel the common factor of y in (xy)/(y).
x=(0)/(y)
0 divided by any number or variable is 0.
x=0
or x=a value yet y=0 because the two multiplied together equals zero.